3. An iron rod 4 m long and 0.5 cm2 in cross-section stretches 1 mm when a mass of 225 kg is hung from its lower end. Compute Young's modulus for the iron.
1
Expert's answer
2020-07-13T11:44:52-0400

The relation between stretch and load is given by the Hooke's law:


"\dfrac{mg}{A} = Y\dfrac{\Delta L}{L}"

where "m = 225kg" is the mass of the load, "g = 9.81m\/s^2" is the constant, "\Delta L = 1mm = 10^{-3}m" is the strech, "L = 4m" is the length of the rod, "A = 0.5cm^2 = 0.5\cdot 10^{-4}m^2" and "Y" is the Young's modulus.

Expressing the Young's modulus, get:


"Y = \dfrac{mgL}{A\Delta L} = \dfrac{225\cdot 9.81\cdot4}{0.5\cdot 10^{-4}\cdot 10^{-3}} = 176.58\cdot 10^9Pa = 176.58GPa"

Answer. 176.58 GPa.